chazz5:

A principal of $4600 is invested at 7.5% interest, compounded annually. How much will the investment be worth after 14 years? Use the calculator provided and round your answer to the nearest dollar.

chazz5:

@aeon

Luigi0210:

Do you at least know the formula you're going to use?

chazz5:

yeah the Exponential growth

Luigi0210:

Exponential growth? Are you sure it's not the Compound Interest formula?

chazz5:

yeah you right

chazz5:

you got a answer?

Luigi0210:

We will define the formula as such: \[\Large A= P(1+\frac{r}{n} )^{nt} \] Where P is the principle amount, r is the interest rate, n is the number of times interest is applied, and t is the time. just plug in the information: \[\Large A = 4600(1+ \frac{0.075}{1} ) ^{1*14} \] So just plug this into a calculator: \[\Large A = 4600(1+0.075)^{14} \]

digitalprincess:

A = 3400(1 + 0.055)11 Can you calculate and answer?